OpenStudy (anonymous):

find the cube roots of the complex number 27(cos 11pi/6+i sin 11pi6)

OpenStudy (anonymous):

take the cube root of 27, get 3 divide the angles by 3

OpenStudy (anonymous):

...

OpenStudy (anonymous):

is that 33pi/0?

OpenStudy (anonymous):

that is not the answer im looking for....

OpenStudy (anonymous):

\[ \frac {33\pi} 6 \]

OpenStudy (anonymous):

my answer is suppose to be 3(cos 11pi/18 + i sin 11pi/18)

OpenStudy (anonymous):

Sorry, I did \[ x^3 \] instead 0f \[x^{\frac 1 3} \]

OpenStudy (anonymous):

and 3(cos 23pi/18 + i sin 23pi/18) and 3(cos 35pi/18 + i sin 35pi/18)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ x = 27 e^{ i \left( \frac {11 \pi}{ 6}+ 2 k \pi \right) }\\ x^{\frac 1 3}= 27^{\frac 1 3 } e^{ i \left( \frac {11 \pi}{ 18}+ 2 k\frac \pi 3 \right) }\\ \]

OpenStudy (anonymous):

Take k=0,1,2 and you get your three answers.

OpenStudy (anonymous):

@lorim9, you got this?

OpenStudy (anonymous):

i think i do now.

OpenStudy (anonymous):

divide \(\frac{11\pi}{6}\) by 3, get \(\frac{11\pi}{18}\) add \(2\pi\) to \(\frac{11\pi}{6}\) get \(\frac{23\pi}{6}\) divide by 3 again get \(\frac{23\pi}{18}\)

OpenStudy (anonymous):

what you and the other person is kinda similar?

OpenStudy (anonymous):

and once more, add \(2\pi\) to \(\frac{23\pi}{6}\) get \(\frac{35}{6}\) divide once more and get \(\frac{35}{18}\) those are your three angles. yes it is exactly the same

OpenStudy (anonymous):

alright thank you guys!